# Codes Correcting Two Bursts of Exactly $b$ Deletions

### v3

@@ Abstract
In this paper, we study codes that correct two bursts of deletions, with each burst having length exactly $b$, where $b>1$. Previously, the best known construction, derived using the syndrome compression technique, achieved at most $7\log n+O\left(\log n/\log\log n\right)$ redundant bits. In this work, we develop new idea to construct $q$-ary codes that achieve redundancy at most $6\log n+O(\log\log n)$ for all $b>1$ and $q\geq 2$. At last, for $b=1$, we provide a new construction of $q$-ary two-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ for all $q>2$.
@@ I Introduction
A subset $\mathcal{C}\subseteq\{0,1,\ldots,q-1\}^{n}$ (where $q\geq 2$) is called a $t$-deletion correcting code, if it has the property that if a codeword $\mathbf{x}\in\mathcal{C}$ is corrupted by deleting $t$ symbols to obtain a subsequence $\mathbf{y}\in\{0,1,\ldots,q-1\}^{n-t}$, then one can recover $\mathbf{x}$ from $\mathbf{y}$. The study of deletion correcting codes has a long history, dating back to at least the 1960s [1]. The seminal work in this field is [2], where by proposing a linear-time decoding algorithm, Levenshtein proved that the binary code (Varshamov-Tenengolts code, or VT code for short) constructed in [3] can combat a single deletion error. In 1984, by leveraging the VT code, Tenengolts constructed a non-binary code (Tenengolts code) that can correct a single deletion [4]. For fixed $t$, $q$ and growing $n$, which is the regime of interest in this paper, the optimal redundancy of a $t$-deletion correcting code $\mathcal{C}$ of length $n$, defined as $\log\left(q^{n}/\left|C\right|\right)$ [^1] All logarithms in this paper are to the base $2$, is asymptotically lower bounded by $t\log n+o(\log n)$ [5] (for $q=2$, the lowere bound is $t\log n+\Omega(1)$ [2]) and upper bounded by $2t\log n-\log\log n+O(1)$ [6]. This implies that codes in [3, 4] have redundancy optimal up to a constant.
@@
Due to applications in DNA-based data storage [7, 8, 9], document exchange [10, 11], multiple-deletion correcting codes with low redundancy have attracted a lot of interest in recent years [10, 12, 11, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. To the best of our knowledge, the best known binary $2$-deletion correcting code with redundancy $4\log n+O(\log\log n)$ was given in [19]. For general $t\geq 3$, the smallest redundancy, which is $(4t-1)\log n+o(\log n)$, was achieved by a construction given in [20]. For non-binary alphabets, Sima et al [15] presented a family of $q$-ary $t$-deletion correcting codes with redundancy $4t\log n+o(\log n)$ by using the syndrome compression technique. The syndrome compression technique was improved in [20] to the so-called syndrome compression technique with pre-coding. A straightforward application of this method will give a $q$-ary $t$-deletion correcting codes with redundancy $(4t-1)\log n+o(\log n)$, which is the best known result in redundancy. When $q>2$ is even and $t=2$, Song and Cai recently constructed a class of $q$-ary $2$-deletion correcting codes with redundancy $5\log(n)+O(\log\log(n))$ [21]. And a following work [22] gave a $q$-ary $2$-deletion correcting code with redundancy $5\log(n)+O(\log\log(n))$ for all $q>2$. When $t=1$, Nguyen et al recently constructed a new $q$-ary single-deletion correcting code with redundancy $\log n+\log q$ [23]. In addition, they showed that there is a linear time encoder with near-optimal redundancy for their code.
@@
If deletions occur at consecutive positions, we call them a burst of deletions. Codes correcting this type of error are of interest due to applications in DNA-based data storage [7, 24], wireless sensor networks and satellite communication devices [25]. A code is called a $b$-burst-deletion correcting code, if it can correct any single burst of exactly $b$ deletions. In 1970, Levenshtein gave a class of binary codes with redundancy at most $\log n+1$ when $b=2$ [26]. For $b\geq 3$, Cheng et al in 2014 constructed a class of binary codes with redundancy $b\log(n/b+1)$ [27], which was later improved to $\log n+(b-1)\log\log n+O(1)$ by Schoeny et al in 2017 [28]. Schoeny’s result was generalized to non-binary alphabets in [29, 30]. The best known redundancy for all $q\geq 2$ is $\log n+O(1)$, which was contributed recently by Sun et al[31]. It was proved in [28, 30] that the redundancy of a $b$-burst-deletion correcting code is at least $\log n+\Omega(1)$. Therefore, the codes in [26] and [31, Theorem 9] have redundancy optimal up to a constant. There are also a lot of works on codes correcting single burst of at most $b$ deletions [26, 28, 32, 33, 21, 34, 35, 31].
@@
In this work, we focus on codes correcting two bursts of deletions, where each burst is of length exactly $b$. To the best of our knowledge, there is no explicit results about such codes. A related result can be found in a work of Sima et al [32, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ where the deletions in each burst need not occur consecutively. Let $t=2$ and $t_{L}=b$. Then their result gives a binary $2$-$b$-burst-deletion correcting code with redundancy at most $8\log n+o(\log n)$. A straightforward application of the syndrome compression technique with pre-coding incurs a code with redundancy at most $7\log n+O(\log n/\log\log n)$, for all $q\geq 2$. This conclusion also holds when $b=1$, i.e., for the case of two-deletion correcting codes. On the other hand, [17, 19] already confirmed that there are two-deletion correcting codes outperforming the one given by the syndrome compression technique. Specifically, [17] gave a binary code with redundancy at most $7\log n+O(1)$ and [19] gave a binary code with redundancy at most $4\log n+O(\log\log n)$. Note that in a burst of deletions, all deletions occur consecutively. Therefore, intuitively, it is reasonable to deem that there is no big difference between two-deletion correcting codes and codes for correcting two $b$-burst-deletions (where $b>1$). This raises a natural question: for $b>1$, is there a construction which leads to codes which are as good as or even better than the one given by the syndrome compression technique. Motivated by this question, in this paper, we investigate new constructions of codes for correcting two $b$-burst-deletions for all $b\geq 2$. Our contributions include:
@@
  • A binary code of length $n$ with redundancy at most $6\log n+(15b-5)\log\log n+O(1)$;
  • A $q$-ary code of length $n$ with redundancy at most $6\log n+(14b\left\lceil\log q\right\rceil+11)\log\log n+O(1)$, for all $q\geq 2$;
@@
Here, it is assumed that $q$ and $b$ are constants with respect to $n$.
The rest of this paper is organized as follows. In Section II, we introduce some necessary definitions and related results. In Section III, we bound above and below the size (or equivalently, the redundancy) of codes correcting two bursts of exactly $b$ deletions. Section IV and Section V deal with binary and non-binary codes, respectively. In Section VI, we show that our idea in Section V can result in a new construction of non-binary two-deletion correcting code. Finally Section VII concludes this paper.
@@ II Preliminaries
In this section, we introduce some necessary terminologies and notations.
@@
Let $\mathcal{C}$ be a subset of $\Sigma_{q}^{n}$ with $\left|\mathcal{C}\right|\geq 2$. Suppose $t\in\{1,2\}$. We call $\mathcal{C}$ a $t$-$b$-burst-deletion correcting code if $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})=\emptyset$ for any two distinct $\mathbf{x},\mathbf{y}\in\mathcal{C}$. In particular, if $b=1$, we call $\mathcal{C}$ a $t$-deletion correcting code.
@@
We call $A\left(\mathbf{x},n,n^{\prime}\right)$ a matrix representation of $\mathbf{x}$. If $n$ and $n^{\prime}$ are clear from the context, we will denote $A(\mathbf{x},n,n^{\prime})$ by $A(\mathbf{x})$.
@@
In this paper, when dealing with $t$-$b$-burst-deletion correcting codes, it is helpful to represent a sequence $\mathbf{x}$ of length $n$ as an array $A(\mathbf{x},n,b)$. To avoid ceiling functions (for example, $\left\lceil n/b\right\rceil$), we always assume $b\mid n$. All results in this paper still hold even if $b\nmid n$, as long as we replace $n/b$ by $\left\lceil n/b\right\rceil$.
The next observation is a straightforward corollary of II.1.
@@ II-A Related Results
Recently, there have been two notable developments in non-binary two-deletion correcting codes with low redundancy. In [21, Theorem 1], the authors presented a $q$-ary two-deletion correcting code with redundancy at most $5\log n+(16\log q+10)\log\log n+o(\log\log n)$, for all even $q>2$. In [22], a $q$-ary two-deletion correcting code was constructed, with redundancy at most $5\log n+10\log\log n+O_{q}(1)$ (where $O_{q}(1)$ denotes a constant depending only on $q$), for all $q>2$. In Section VI, we present a new construction of $q$-ary two-deletion correcting codes with redundancy at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$, for all $q>2$.
Regarding codes that can correct two $b$-burst-deletions (where $b>1$), there are, to our knowledge, no explicit results available. A related result can be found in the work of Sima et al [32, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ with the deletions in each burst not necessarily occurring consecutively. For $t=2$ and $t_{L}=b$, their result provides a binary $2$-$b$-burst-deletion correcting code with redundancy at most $8\log n+o(\log n)$. Their result was derived using the syndrome compression technique, which was later extended to syndrome compression with pre-coding in [20]. We will apply this extended technique to provide a $q$-ary code with redundancy at most $7\log n+O\left(\log n/\log\log n\right)$, for all $q\geq 2$.
@@ 1. $(\textup{{P}}1)$
[31, Theorem 9, $t=b,s=0$] For all $q\geq 2$ and $n>b$, there is a function $\phi:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{\log n+O_{q,b}(1)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\phi(\mathbf{x})$ and $\mathbf{y}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$. Here, $O_{q,b}(1)$ is a constant dependent only on $q$ and $b$.
This lemma gives a $1$-$b$-burst-deletion correcting code $\mathcal{E}$ with redundancy $\log n+O_{q,b}(1)$. Since $\mathcal{E}$ can correct single $b$-burst-deletion, by a simple counting, we can see that $\left|\mathcal{N}_{\mathcal{E}}(\mathbf{x})\right|\leq q^{b}n^{3}$. So we let $N=q^{b}n^{3}$. Now the redundancy of $\mathcal{E}^{\prime}$ is at most $7\log n+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)$. To conclude our discussion, it remains to find an $f$ satisfying (P1) in Lemma II.1 such that $R(n)=O(\log n)$, which is given in Lemma II.5. The proof of Lemma II.5 is based on Lemmas II.3 and II.4.
@@
Before proceeding to subsequent sections, we introduce two useful lemmas. The first one is a well-known result from number theory, which will be frequently used. The second one will be used in Sections IV-A and V-A.
@@
Let $N>2$ be an integer. Suppose $\mathbf{x}\in[0,N-1]^{n}$, where $n>4$. Let ???? denote an unknown symbol (not in $[0,N-1]$). If $\mathbf{y}\in\left([0,N-1]\cup\{?\}\right)^{n}$ be such that $y_{i}=y_{i+1}=y_{j}=y_{j+1}=?$ and $y_{k}=x_{k}$ for all $k\notin\{i,i+1,j,j+1\}$, we say that $\mathbf{y}$ is obtained from $\mathbf{x}$ by two bursts of erasures (of length two). For our purpose in Sections IV-A and V-A, assume $j\geq i+2$. For a sequence $\mathbf{x}$ over the alphabet $[0,N-1]$, denote $A(\mathbf{x})=A(\mathbf{x},\left|\mathbf{x}\right|,2)$ and $\textup{Syn}\left(\mathbf{x}\right)\triangleq\sum_{i=1}^{n}ix_{i}$.
@@
Let $Q\geq N$ be the smallest prime. For all $0\leq a_{1},a_{2},b_{1},b_{2}<Q$ and $0\leq c<nN^{2}$, define $\mathcal{C}$ to be set of all sequences $\mathbf{x}\in[0,N-1]^{n}$ that satisfies:
@@
    $\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{1},j\equiv a_{1}\pmod{Q}$, $\sum_{j=1}^{\left\lceil n/2\right\rceil}\left(A(\mathbf{x})_{1},j\right)^{2}\equiv a_{2}\pmod{Q}$, $\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2},j\equiv b_{1}\pmod{Q}$,
    $\sum_{j=1}^{\left\lceil n/2\right\rceil}\left(A(\mathbf{x})_{2},j\right)^{2}\equiv b_{2}\pmod{Q}$;
@@
    $W(\mathbf{x})\equiv c\pmod{nN^{2}}$, where $W(\mathbf{x})=\textup{Syn}\left(A(\mathbf{x})_{1}\right)+2N\cdot\textup{Syn}\left(A(\mathbf{x})_{2}\right)$.
Then the code $\mathcal{C}$ can correct two bursts of erasures. In particular, there is a function $\mathcal{E}_{1}:[0,N-1]^{n}\rightarrow\Sigma_{2}^{\log n+6\log N+4}$, computable in linear time, such that for any $\mathbf{x}\in[0,N-1]^{n}$, given $\mathcal{E}_{1}(\mathbf{x})$, we can efficiently and uniquely recover $\mathbf{x}$ from $\mathbf{y}$, where $\mathbf{y}$ is any given sequence obtained from $\mathbf{x}$ by two bursts of erasures.
@@
$\mathcal{E}_{1}(\mathbf{x})\triangleq\left(\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{1},j\pmod{Q},\sum_{j=1}^{\left\lceil n/2\right\rceil}\left(A(\mathbf{x})_{1},j\right)^{2}\pmod{Q},\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2},j\pmod{Q},\right.$
$\left.\sum_{j=1}^{\left\lceil n/2\right\rceil}\left(A(\mathbf{x})_{2},j\right)^{2}\pmod{Q},W(\mathbf{x})\pmod{nN^{2}}\right).$
@@
Here, we view $\mathcal{E}(\mathbf{x})$ as a binary vector. By Lemma II.6, we have $N\leq Q<2N$. Now it is easy to see that the length of $\mathcal{E}(\mathbf{x})$ is at most $\log n+6\log N+4$. It remains to prove the correctness of $\mathcal{C}$.

### v4

@@ Abstract
In this paper, we investigate codes designed to correct two bursts of deletions, where each burst has a length of exactly $b$, where $b>1$. The previous best construction, achieved through the syndrome compression technique, had a redundancy of at most $7\log n+O\left(\log n/\log\log n\right)$ bits. In contrast, our work introduces a novel approach for constructing $q$-ary codes that attain a redundancy of at most $5\log n+O(\log\log n)$ bits for all $b>1$ and $q\geq 2$. Additionally, for the case where $b=1$, we present a new construction of $q$-ary two-deletion correcting codes with a redundancy of $5\log n+O(\log\log n)$ bits, for all $q>2$.
@@ I Introduction
A subset $\mathcal{C}\subseteq\{0,1,\ldots,q-1\}^{n}$ (where $q\geq 2$) is called a $t$-deletion correcting code, if it has the property that if a codeword $\mathbf{x}\in\mathcal{C}$ is corrupted by deleting $t$ symbols to obtain a subsequence $\mathbf{y}\in\{0,1,\ldots,q-1\}^{n-t}$, then one can recover $\mathbf{x}$ from $\mathbf{y}$. The study of deletion correcting codes has a long history, dating back to at least the 1960s [1]. The seminal work in this field is [2], whereby proposing a linear-time decoding algorithm, Levenshtein proved that the binary code (Varshamov-Tenengolts code, or VT code for short) constructed in [3] can combat a single deletion error. In 1984, by leveraging the VT code, Tenengolts constructed a non-binary code (Tenengolts code) that can correct a single deletion [4]. For fixed $t$, $q$ and growing $n$, which is the regime of interest in this paper, the optimal redundancy of a $t$-deletion correcting code $\mathcal{C}$ of length $n$, defined as $\log\left(q^{n}/\left|C\right|\right)$ [^1] All logarithms in this paper are to the base $2$, is asymptotically lower bounded by $t\log n+o(\log n)$ [5] (for $q=2$, the lower bound is $t\log n+\Omega(1)$ [2]) and upper bounded by $2t\log n-\log\log n+O(1)$ [6]. This implies that the VT code in [3] and the Tenengolts code [4] have redundancy optimal up to a constant.
@@
Due to applications in DNA-based data storage [7, 8, 9], document exchange [10, 11], multiple-deletion correcting codes with low redundancy have attracted a lot of interest in recent years [10, 12, 11, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. To the best of our knowledge, the best known binary $2$-deletion correcting code with redundancy $4\log n+O(\log\log n)$ was given in [19, 24]. For general $t\geq 3$, the smallest redundancy, which is $(4t-1)\log n+o(\log n)$, was achieved by a construction given in [20]. For non-binary alphabets, Sima et al [15] presented a family of $q$-ary $t$-deletion correcting codes with redundancy $4t\log n+o(\log n)$ by using the syndrome compression technique. The syndrome compression technique was improved in [20] to the so-called syndrome compression technique with pre-coding. A straightforward application of this method will give a $q$-ary $t$-deletion correcting codes with redundancy $(4t-1)\log n+o(\log n)$, which is the best-known result in redundancy. When $q>2$ is even and $t=2$, Song and Cai recently constructed a class of $q$-ary $2$-deletion correcting codes with redundancy $5\log n+O(\log\log n)$ [21]. And in a following work [22] the authors presented a $q$-ary $2$-deletion correcting code with redundancy $5\log n+O(\log\log n)$ for all $q>2$. When $t=1$, Nguyen et al recently constructed a new $q$-ary single-deletion correcting code with redundancy $\log n+\log q$ [23]. In addition, they showed that there is a linear time encoder with near-optimal redundancy for their code.
@@
If deletions occur at consecutive positions, we call them a burst of deletions. Codes correcting this type of error are of interest due to applications in DNA-based data storage [7, 25], wireless sensor networks, and satellite communication devices [26]. A code is called a $b$-burst-deletion correcting code, if it can correct any single burst of exactly $b$ deletions. In 1970, Levenshtein presented a class of binary codes with redundancy at most $\log n+1$ when $b=2$ [27]. For $b\geq 3$, Cheng et al in 2014 constructed a class of binary codes with redundancy $b\log(n/b+1)$ [28], which was later improved to $\log n+(b-1)\log\log n+O(1)$ by Schoeny et al in 2017 [29]. Schoeny’s result was generalized to non-binary alphabets in [30, 31]. The best known redundancy for all $q\geq 2$ is $\log n+O(1)$, which was contributed recently by Sun et al[32]. It was proved in [29, 31] that the redundancy of a $b$-burst-deletion correcting code is at least $\log n+\Omega(1)$. Therefore, the codes in [27] and [32, Theorem 9] have redundancy optimal up to a constant. There are also a lot of works on codes correcting single burst of at most $b$ deletions [27, 29, 33, 34, 21, 35, 36, 32].
@@
In this work, we focus on codes correcting two bursts of deletions, where each burst is of length exactly $b$. We call such codes $2$-$b$-burst correcting codes. To the best of our knowledge, there are no explicit results about such codes. A related result can be found in a work of Sima et al [33, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ where the deletions in each burst need not occur consecutively. Let $t=2$ and $t_{L}=b$. Then their result gives a binary $2$-$b$-burst-deletion correcting code with redundancy at most $8\log n+o(\log n)$. A straightforward application of the syndrome compression technique with pre-coding incurs a code with redundancy at most $7\log n+O(\log n/\log\log n)$, for all $q\geq 2$. This conclusion also holds when $b=1$, i.e., for the case of two-deletion correcting codes. On the other hand, [17, 19, 21, 22, 24] already confirmed that there are two-deletion correcting codes outperforming the one given by the syndrome compression technique. Specifically, [17] and [19, 24] presented binary codes with redundancy at most $7\log n+O(1)$ and $4\log n+O(\log\log n)$, respectively. For non-binary codes, the best-known redundancy is $5\log n+O(\log\log n)$ [21, 22]. Note that in a burst of deletions, all deletions occur consecutively. Therefore, it is reasonable to deem by intuition that there is no big difference between two-deletion correcting codes and $2$-$b$-burst-deletion correcting codes (where $b>1$). This raises a natural question: for $b>1$, is there a construction that leads to codes that are as good as, or even better than, the one given by the syndrome compression technique? Motivated by this question, in this paper, we investigate new constructions of codes for correcting two $b$-burst-deletions for all $b\geq 2$. Our contributions include:
@@
  • A binary $2$-$b$-burst-deletion correcting code of length $n$ with redundancy at most $5\log n+14b\log\log n+O(1)$, for all $b>1$;
  • A $q$-ary $2$-$b$-burst-deletion correcting code of length $n$ with redundancy at most $5\log n+(14b\left\lceil\log q\right\rceil+14)\log\log n+O(1)$, for all $q>2$ and $b>1$;
@@
Here, it is assumed that $q$ and $b$ are constants with respect to $n$. Therefore, our results show that for $2$-$b$-burst-deletion correcting codes, we can do almost as well as two-deletion correcting codes.
The rest of this paper is organized as follows. In Section II, we introduce some necessary definitions and related results. In Section III, we bound above and below the size (or equivalently, the redundancy) of codes correcting two bursts of exactly $b$ deletions. Section IV deals with codes for correcting two $b$-burst-deletions. In Section V, we give a new construction of non-binary two-deletion correcting code. Finally Section VI concludes this paper.
@@ II Preliminaries
In this section, we first introduce some necessary definitions, auxiliary conclusions and related results.
@@
Let $\mathcal{C}$ be a subset of $\Sigma_{q}^{n}$ with $\left|\mathcal{C}\right|\geq 2$. Suppose $t\in\{1,2\}$. We call $\mathcal{C}$ a $t$-$b$-burst-deletion correcting code if $\mathcal{B}_{t}^{b}(\mathbf{x})\cap\mathcal{B}_{t}^{b}(\mathbf{y})=\emptyset$ for any two distinct $\mathbf{x},\mathbf{y}\in\mathcal{C}$. In particular, if $b=1$, we call $\mathcal{C}$ a $t$-deletion correcting code.
@@
We call $A\left(\mathbf{x},n,n^{\prime}\right)$ a matrix (or array) representation of $\mathbf{x}$. If $n$ and $n^{\prime}$ are clear from the context, we will denote $A(\mathbf{x},n,n^{\prime})$ by $A(\mathbf{x})$.
@@
In this paper, when dealing with $t$-$b$-burst-deletion correcting codes, it is helpful to represent a sequence $\mathbf{x}$ of length $n$ as matrix $A(\mathbf{x},n,b)$. To avoid ceiling functions (for example, $\left\lceil n/b\right\rceil$), we always assume $b\mid n$. All results in this paper still hold even if $b\nmid n$, as long as we replace $n/b$ by $\left\lceil n/b\right\rceil$. Throughout this paper, for a matrix $A$, denote by $A_{i}$ and $A_{i,j}$ the $i$-th row of $A$ and the entry in the $i$-th row and $j$-th column, respectively.
The next observation is a straightforward result of II.1.
@@ II-A Related Results
Recently, there have been two notable developments in non-binary two-deletion correcting codes with low redundancy. In [21, Theorem 1], the authors presented a $q$-ary two-deletion correcting code with redundancy at most $5\log n+(16\log q+10)\log\log n+o(\log\log n)$, for all even $q>2$. In [22], a $q$-ary two-deletion correcting code was constructed, with redundancy at most $5\log n+10\log\log n+O_{q}(1)$ (where $O_{q}(1)$ denotes a constant depending only on $q$), for all $q>2$. In Section V, we present a new construction of $q$-ary two-deletion correcting codes with redundancy at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$, for all $q>2$.
Regarding codes that can correct two $b$-burst-deletions (where $b>1$), there are, to our knowledge, no explicit results available. A related result can be found in the work of Sima et al [33, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ with the deletions in each burst not necessarily occurring consecutively. For $t=2$ and $t_{L}=b$, their result provides a binary $2$-$b$-burst-deletion correcting code with redundancy at most $8\log n+o(\log n)$. Their result was derived using the syndrome compression technique, which was later extended to syndrome compression with pre-coding in [20]. We will apply this extended technique to provide a $q$-ary code with redundancy at most $7\log n+O\left(\log n/\log\log n\right)$, for all $q\geq 2$.
@@ 1. $(\textup{{P}}1)$
[32, Theorem 9, $t=b,s=0$] For all $q\geq 2$ and $n>b$, there is a function $\phi:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{\log n+O_{q,b}(1)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\phi(\mathbf{x})$ and $\mathbf{y}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$. Here, $O_{q,b}(1)$ is a constant dependent only on $q$ and $b$.
This lemma gives a $1$-$b$-burst-deletion correcting code $\mathcal{E}$ with redundancy $\log n+O_{q,b}(1)$. Since $\mathcal{E}$ can correct single $b$-burst-deletion, by a simple counting, we can see that $\left|\mathcal{N}_{\mathcal{E}}(\mathbf{x})\right|\leq q^{b}n^{3}$. In fact, each codeword in $\mathcal{N}_{\mathcal{E}}(\mathbf{x})$ can be obtained in the following three steps:
  1. 1)
    Delete two substrings of length $b$ from $\mathbf{x}$, resulting in $\mathbf{z}^{(1)}$. There are at most $n^{2}$ possibilities for $\mathbf{z}^{(1)}$.
  2. 2)
    For each $\mathbf{z}^{(1)}$, insert a sequence of length $b$ into $\mathbf{z}^{(1)}$ and get a sequence $\mathbf{z}^{(2)}\in\Sigma_{q}^{n-b}$. For each $\mathbf{z}^{(1)}$, there are at most $q^{b}n$ possibilities for $\mathbf{z}^{(2)}$.
  3. 3)
    Insert a sequence of length $b$ into $\mathbf{z}^{(2)}$ to get a sequence $\mathbf{y}\in\Sigma_{q}^{n}$. Since $\mathcal{E}$ is a $1$-$b$-burst-deletion correcting code, for each $\mathbf{z}^{(2)}$, there is at most one $\mathbf{y}$ which is in $\mathcal{N}_{\mathcal{E}}(\mathbf{x})$.
Therefore, we can let $N=q^{b}n^{3}$. Now the redundancy of $\mathcal{E}^{\prime}$ is at most $7\log n+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)$. To conclude our discussion, it remains to find an $f$ satisfying (P1) in Lemma II.1 such that $R(n)=O(\log n)$, which is given in Lemma II.5. The proof of Lemma II.5 is based on Lemmas II.3 and II.4.
@@
This result can be extended to arbitrary finite alphabets in the following way.
@@
Before proceeding to subsequent sections, we introduce two useful lemmas. The first one is a well-known result from number theory. The second one will be used in Section IV-B.
@@
Let $N\geq 2$ be an integer. Suppose $\mathbf{x}\in[0,N-1]^{n}$, where $n>4$. Let ???? denote an unknown symbol (not in $[0,N-1]$). If $\mathbf{y}\in\left([0,N-1]\cup\{?\}\right)^{n}$ be such that $y_{i}=y_{i+1}=y_{j}=y_{j+1}=?$ and $y_{k}=x_{k}$ for all $k\notin\{i,i+1,j,j+1\}$, we say that $\mathbf{y}$ is obtained from $\mathbf{x}$ by two bursts of erasures (of length two). For our purpose in Section IV-B, assume $j\geq i+2$. For a sequence $\mathbf{x}$ over the alphabet $[0,N-1]$, denote $A(\mathbf{x})=A(\mathbf{x},\left|\mathbf{x}\right|,2)$ and $\textup{Syn}\left(\mathbf{x}\right)\triangleq\sum_{i=1}^{n}ix_{i}$.
@@
For any $0\leq a_{1},a_{2}<2N$ and $0\leq b<nN^{2}$, define $\mathcal{C}$ to be set of all sequences $\mathbf{x}\in[0,N-1]^{n}$ that satisfies:
@@
    $\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{1,j}\equiv a_{1}\pmod{2N}$, $\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2,j}\equiv a_{2}\pmod{2N}$;
@@
    $W(\mathbf{x})\equiv b\pmod{nN^{2}}$, where $W(\mathbf{x})=\textup{Syn}\left(A(\mathbf{x})_{1}\right)+(2N-1)\cdot\textup{Syn}\left(A(\mathbf{x})_{2}\right)$.
Then the code $\mathcal{C}$ can correct two bursts (of length two) of erasures. In particular, there is a function $\varphi:[0,N-1]^{n}\rightarrow\Sigma_{2}^{\log n+4\log N+2}$, efficiently computable, such that for any $\mathbf{x}\in[0,N-1]^{n}$, given $\varphi(\mathbf{x})$, we can efficiently and uniquely recover $\mathbf{x}$ from $\mathbf{y}$, where $\mathbf{y}$ is any given sequence obtained from $\mathbf{x}$ by two bursts of erasures.
@@
$\varphi(\mathbf{x})\triangleq\left(\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{1,j}\pmod{2N},\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2,j}\pmod{2N},W(\mathbf{x})\pmod{nN^{2}}\right).$
@@
Here, we view $\varphi(\mathbf{x})$ as a binary vector. There are at most $(2N)^{2}\cdot(nN^{2})$ values of $\varphi(\mathbf{x})$. As a result, the length of $\varphi(\mathbf{x})$, when viewed as a binary vector, is at most $\log n+4\log N+2$. It remains to prove the correctness of $\mathcal{C}$.